Ideal Co. Ltd is a local company which produces medical mask. Due to the outbreak of Coronavirus in Hong Kong, the CEO considers buying a new set of machines to enlarge the production scale. TWO models of the same budget ($200,000) are available and the information is shown below:
Model A Model B
Resell value after ONE year $115,000 $135,000

Annual repair and maintenance fee $25,000 $20,000
Annual electricity consumption$30,000 $75,000
Which set of machines will you suggest the Company to purchase? Explain briefly with relevant data and working.
1
Expert's answer
2020-05-19T10:08:26-0400

If we don't count the discount rate, then the total cost of model A is:

TCa = 200,000 - 115,000 + 25,000 + 30,000 = 140,000.

The total cost of model A is:

TCb = 200,000 - 135,000 + 20,000 + 75,000 = 160,000.

So, model A is better.